OpenStudy (anonymous):

Let f′′(x)=24x2+6, f(1)=10, f′(1)=4 Then f(x)=?

OpenStudy (anonymous):

integrate f double prime, and then use the fact that f prime=4 to find the constant, then integrate again to find f(x)

OpenStudy (anonymous):

*that f prime at 1 =4, thats key

OpenStudy (lgbasallote):

do you know how to integrate 24x^2 + 6?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

\(\int (24x^2 + 6)dx = \int 24x^2 dx + \int 6dx = 24\int x^2dx + 6\int dx\) now can you integrate it?

OpenStudy (anonymous):

ohhhh

OpenStudy (lgbasallote):

tell me when you have found the integral :)